Genuine Spice Inc. began operations on January 1 of the current year. The company produces eight- ounce bottles of hand and body lotion called Eternal Beauty. The lotion is sold wholesale in 12-bottle cases for $100 per case. There is a selling commission of $20 per case. The January direct materials, direct labor, and factory overhead costs are as follows:

DIRECT MATERIALS
Cost Behavior Units per Case Cost per Unit Cost per Case
Cream base Variable 100 oz. $0.02 $ 2.00
Natural oils Variable 30 oz. 0.30 9.00
Bottle (8-oz.) Variable 12 bottles 0.50 6.00
$17.00
DIRECT LABOR
Department Cost Behavior Time per Case Labor Rate per Hour Cost per Case
Mixing Variable 20 min. $18.00 $6.00
Filling Variable 5 14.40 1.20
25 min. $7.20
FACTORY OVERHEAD
Cost Behavior Total Cost
Utilities Mixed $600
Facility lease Fixed 14,000
Equipment depreciation Fixed 4,300
Supplies Fixed 660
$19,560

Part A—Break-Even Analysis

The management of Genuine Spice Inc. wants to determine the number of cases required to break even per month. The utilities cost, which is part of factory overhead, is a mixed cost. The following information was gathered from the first six months of operation regarding this cost:

Case Production

Utility Total Cost

January 500 $600
February 800 660
March 1,200 740
April 1,100 720
May 950 690
June 1,025 705
Required-Part A:
1. Determine the fixed and variable portion of the utility cost using the high-low method.
2. Determine the contribution margin per case.
3. Determine the fixed costs per month, including the utility fixed cost from part (1).
4. Determine the break-even number of cases per month.

Part B—August Budgets

During July of the current year, the management of Genuine Spice Inc. asked the controller to prepare August manufacturing and income statement budgets. Demand was expected to be 1,500 cases at $100 per case for August. Inventory planning information is provided as follows:

Finished Goods Inventory:

Cases

Cost

Estimated finished goods inventory, August 1 300 $12,000
Desired finished goods inventory, August 31 175 7,000

Materials Inventory:

Cream Base

Oils

Bottles

(oz.)

(oz.)

(bottles)

Estimated materials inventory, August 1 250 290 600
Desired materials inventory, August 31 1,000 360 240

There was negligible work in process inventory assumed for either the beginning or end of the month; thus, none was assumed. In addition, there was no change in the cost per unit or estimated units per case operating data from January.

Required-Part B:
5. Prepare the August production budget.*
6. Prepare the August direct materials purchases budget.*
7. Prepare the August direct labor cost budget. Round the hours required for production to the nearest hour.*
8. Prepare the August factory overhead cost budget. If an amount box does not require an entry, leave it blank. (Entries of zero (0) will be cleared automatically by CNOW.)*
9. Prepare the August budgeted income statement, including selling expenses. NOTE: Because you are not required to prepare a cost of goods sold budget, the cost of goods sold calculations will be part of the budgeted income statement.*
*Enter all amounts as positive numbers.

Part C—August Variance Analysis

During September of the current year, the controller was asked to perform variance analyses for August. The January operating data provided the standard prices, rates, times, and quantities per case. There were 1,500 actual cases produced during August, which was 250 more cases than planned at the beginning of the month. Actual data for August were as follows:

Actual Direct Materials

Price per Unit

Quantity per Case

Cream base $0.016 per oz. 102 oz.
Natural oils $0.32 per oz. 31 oz.
Bottle (8-oz.) $0.42 per bottle 12.5 bottles

Actual Direct

Actual Direct Labor

Labor Rate

Time per Case

Mixing $18.20 19.50 min.
Filling 14.00 5.60 min.
Actual variable overhead $305.00
Normal volume 1,600 cases

The prices of the materials were different from standard due to fluctuations in market prices. The standard quantity of materials used per case was an ideal standard. The Mixing Department used a higher grade labor classification during the month, thus causing the actual labor rate to exceed standard. The Filling Department used a lower grade labor classification during the month, thus causing the actual labor rate to be less than standard

Required-Part C:
10. Determine and interpret the direct materials price and quantity variances for the three materials.
11. Determine and interpret the direct labor rate and time variances for the two departments. Round hours to the nearest tenth of an hour.
12. Determine and interpret the factory overhead controllable variance.
13. Determine and interpret the factory overhead volume variance.
14. Why are the standard direct labor and direct materials costs in the calculations for parts (10) and (11) based on the actual 1,500-case production volume rather than the planned 1,375 cases of production used in the budgets for parts (6) and (7)?
Answer #1

Dear Student,

As per the HOMEWORKLIB POLICY, only the first four questions should be answered. kindly take note of it.

Part 1

Variable cost per unit = (Total cost at highest production – total cost at lowest production) / (highest production – lowest production) = (740-600)/(1200-500) = $0.20 per case

At lowest point (500 cases)

Total Cost = Variable Cost + Fixed Cost

600 = (0.20*500) + Fixed cost

Fixed cost = 600-100 = $500

Part 2

Selling price per case

100

Less variable cost per case:

Direct materials

17

Direct labor

7.20

Utilities

0.20

Selling expenses

20

Total variable cost per case

44.40

Contribution margin per case

$55.60

Part 3

utilities

500

Facility lease

14000

Equipment depreciation

4300

Supplies

660

Total fixed costs

$19460

Part 4

Break even sales (in units) = fixed costs / contribution margin per unit = 19460/55.60 = 350 cases